A doctors office provides water coolers with cone- shaped paper cups. A cup has a diameter of 2.5 in., and the distance from the tip of the cone to the lip is 4in. What is the height of the cup? What is the volume of the cup? Round you answers to the nearest tenth.


1
Expert's answer
2022-02-11T12:48:26-0500
"r=2.5\ in./2=1.25\ in."

"h=\sqrt{l^2-r^2}=\sqrt{(4\ in.)^2-(1.25\ in.)^2}\approx3.8\ in."

The height of the cup is "3.8\ in."


"V=\dfrac{\pi r^2 h}{3}=\dfrac{\pi (1.25\ in.)^2\sqrt{(4\ in.)^2-(1.25\ in.)^2} }{3}"

"\approx6.2\ in.^3"

The volume of the cup is 6.2 cubic inches.